### v1

@@ Abstract
Recent trends within computational and data sciences show an increasing recognition and adoption of computational workflows as tools for productivity, reproducibility, and democratized access to platforms and processing know-how. As digital objects to be shared, discovered, and reused, computational workflows benefit from the FAIR principles, which stand for Findable, Accessible, Interoperable, and Reusable. The Workflow Community Initiative FAIR Workflows Working Group (WCI-FW), a global and open community of researchers and developers working with computational workflows across disciplines and domains, has systematically addressed the application of both FAIR data and software principles to computational workflows. We present our recommendations with commentary that reflects our discussions and justifies our choices and adaptations. Like the software and data principles on which they are based, these are offered to workflow users and authors, workflow management system developers, and providers of workflow services as guide rails for adoption and fodder for discussion. Workflows are becoming more prevalent as documented, automated instruments for data analysis, data collection, AI-based predictions, and simulations. The FAIR recommendations for workflows that we propose in this paper will maximize their value as research assets and facilitate their adoption by the wider community.
@@
1 Computational workflows and why FAIR matters
Scale-ups in research data and the rise of data-driven science have spurred the research community to find solutions for managing and automating complex computational processes, ensuring efficiency, reproducibility, scalability, collaboration, and quality-assured transparency. Computational workflows are a special kind of software specifically targeted at handling multi-step, multi-code data pipelines, data analyses, and other data-handling operations, especially through the efficient use of computational resources to transform data inputs into desired outputs.
@@ 1 Computational workflows and why FAIR matters
At one end of the scale, the workflow language might be a simple script (e.g. Bash, Python, R) or enumerated stages in an electronic research notebook (e.g. Jupyter, RStudio, Apache Zeppelin) with a set of instructions that use inputs and outputs piping the results together. At the other end, workflows might employ a workflow management system (WMS) such as Nextflow [2], Galaxy [3], Snakemake [4], and Parsl [5]. Using a WMS can provide a number of benefits including abstraction, scaling, automation, reproducibility [6], and provenance [7]. A fully featured WMS typically facilitates error handling and restarting, automatic restarting and data staging, provenance recording, the handling of large datasets and complex computations, task and resource allocation, and distributed task execution to scale over local workstations, cloud resources, or high-performance computing clusters. WMSes and modular containerized (e.g. Docker, Singularity) software components aid portability and reproducibility. In practice, researchers use mixtures (e.g. chaining scripts from a WMS that in turn is launched from a notebook).
@@
A workflow is the formal specification of the data flow and execution control between executable components, and the expected datasets and parameter files. Its complexity can be composed, and its overall process can be abstracted away from its execution. A workflow run is the instantiation of the workflow with inputs (parameters files, input datasets) and outputs (output data, the provenance execution log and lineage of data products).
@@
The use of workflows has accelerated in the past few years. The existence of more than 350 different workflow management systems of varying maturity evidences the increasing popularity of workflows [8]. Workflows automate repetitive and time-consuming tasks, saving time and computational resources and ensuring that computational experiments can be replicated. They provide a documented record of the computational process which helps in understanding, reviewing, and auditing the workflow. As scientific activity often includes the exploration of analysis variance, modifying workflows to understand effects and changes on data products is simpler when those workflows are clearly described and comparable. Researchers can reuse workflows to re-analyze and adapt to incorporate new methods, tools, or data sources. Sharing and reusing workflows supports team collaboration and standardizes processes and analysis methods.
The Findable, Accessible, Interoperable, and Reusable (FAIR) principles aim to maximize the value and impact of scientific digital objects. Such objects cannot be reused if they cannot be found, accessed, or understood; they cannot be combined if they are not interoperable. The FAIR guidelines for scientific data [9] emphasize making data findable through metadata provision, unique and persistent identifier assignment, ensuring accessibility through open access or controlled access mechanisms, promoting interoperability through standardized formats and metadata schemas, and enabling reusability through clear licensing and documentation. The FAIR principles for research software (FAIR4RS) [10] recommend making software findable through repositories and registries, ensuring accessibility through open licensing and documentation, promoting interoperability through standardization and compatibility with other tools, and enabling reusability through versioning and clear documentation of functionality and dependencies. These principles have made a significant impact on policy and publishing, and they have been taken up by the public and private sector as well as spawning an industry of projects, research programs, and commercial businesses.
2 FAIR Principles applied to Computational Workflows
@@
  • Do we need to tailor software principles to the specific characteristics of workflows? As workflows are executables the principles of software should apply. Nevertheless, the characteristics of the heterogeneous components that make up workflows and the expected user behavior of variant reuse, modification, and portability led us to some customization of the principles. Technical information is needed that details each step’s inputs, outputs, dependencies, and computational requirements as well as configuration files, lists of software dependencies, and other information about the operational context. Missing context is one of the main difficulties faced when porting across computing platforms; applying FAIR to workflows requires fully describing the context necessary for executing the workflows as software.
@@
F1. A workflow is assigned a globally unique and persistent identifier. D-F1
S-F1
@@
F2. A workflow and its components are described with rich metadata. D-F2
S-F2
F3. Metadata clearly and explicitly include the identifier of the workflow, and workflow versions, that they describe. D-F3
S-F3
F4. Metadata and workflow are registered or indexed in a searchable FAIR resource. D-F4
S-F4
A1. Workflow and its components are retrievable by their identifiers using a standardized communications protocol. D-A1
S-A1
A1.1. The protocol is open, free, and universally implementable. D-A1.1
S-A1.1
A1.2. The protocol allows for an authentication and authorization procedure, when necessary. D-A1.2
S-A1.2
A2. Metadata are accessible, even when the workflow is no longer available. D-A2
S-A2
I1. Workflow and its metadata (including workflow run provenance) use a formal, accessible, shared, transparent, and broadly applicable language for knowledge representation. D-I1
S-R1.2
@@
I3. Workflow is specified in a way that allows its components to read, write, and exchange data (including intermediate data), in a way that meets domain-relevant standards. D-R3
S-I1
I4. Workflow and its metadata (including workflow run provenance) include qualified references to other objects and the workflow’s components. D-I3
S-I2
S-R1.2
R1. Workflow is described with a plurality of accurate and relevant attributes. D-R1
S-R1
R1.1. Workflow is released with a clear and accessible license. D-R1.1
S-R1.1
R1.2. Components of the workflow representing levels of granularity are given clear and accessible licenses. D-R1.1
S-R1.1
R1.3. Workflow is associated with detailed provenance of the workflow and of the products of the workflow. D-R1.2
S-R1.2
R2. Workflow includes qualified references to other workflows. D-I3
S-R2
R3. Workflow meets domain-relevant community standards. D-R1.3
S-R3
@@ 2.1 Findability
Rich metadata for the overall workflows and their components (including inputs and outputs) (F2, F3) should use a formal, accessible language for knowledge representation, such as schema.org [^2] https://schema.org, and standardized metadata such as the Bioschemas schema.org profile for Computational Workflows [17] (used for workflow components in Workflow Run Crate profile [18]) and CodeMeta [19]. Other metadata forms for workflows include canonical “Rosetta” languages (e.g. CWL [20], WDL [21]) that abstract from the native workflow language for greater portability and comparability. Such metadata are used for registration and search in dedicated and trusted workflow registries such as WorkflowHub [22] and Dockstore (F4) [23].
2.2 Accessibility
@@
2.3 Interoperability
The data and software principles are combined and blur with principles from Reusability, as we tease apart the FAIR requirements for a workflow, its components (both executable and data), and an instantiated and executed run of the workflow. For example, principle I1 applies to describing the static workflow structure of the specification and capturing dynamic aspects of the run such as execution details and intermediate outputs. The workflow specification plays a role in data in I1, I2, and I4 adapting data principles, as well as the role of metadata of the software, co-opting software principles.
@@ 2.4 Reusability
While discussing findability in Section 2.1, we also highlighted the reusability of workflows as complex compositions intended to be modified, built upon, and incorporated into other workflows, as well as their reusability as runnable executions with modified datasets and parameter files [27]. The data and software principles are adapted to emphasize the composite nature of workflows (both executable and data components) for licensing [28]. Workflows are composed of many parts, potentially from different authors with different intents for how their artifacts may be used; these intents must be clearly stated by the inclusion of licenses that cover all parts of the workflow and sub-parts of the workflows that can be reused as sub-workflows [11].
@@
3 Examples
Here, we provide concrete examples with detailed descriptions for applying the FAIR principles to both workflow specifications and workflow runs.
3.1 Example workflow specification
The Protein MD Setup workflow sets up a protein molecular dynamic simulation that returns a protein structure and simulated 3D trajectories [29]. It is registered in WorkflowHub [^3] https://doi.org/10.48546/workflowhub.workflow.29.3, where it has been assigned a Digital Object Identifier (DOI), which is globally unique (F1). Different versions of this workflow can be identified using unique digital object identifiers (F1.2). The workflow is described using (rich) metadata (F2) that includes the identifier of the workflow and its unique version (F3). The workflow and associated metadata are registered (F4). Moreover, components of the workflow have their own individual persistent identifiers as the workflow is composed of sub-workflow BioBB building blocks from the BioBB Library [^4] https://workflowhub.eu/projects/11 (F1.2). The workflow and its components can be retrieved and downloaded using a standardized HTTPS protocol (A1, A1.1, A1.2). The metadata associated with the workflow and its components is independent from the GitHub code repository and are accessible on WorkflowHub even if the workflow (or its components) become inaccessible (A2).
@@ 3.1 Example workflow specification
The workflow is released under Apache License 2.0 software license (R1.1) and WorkflowHub’s RO-Crate references that the embedded CWL code is archived from the corresponding GitHub source code repository [^5] https://github.com/bioexcel/biobb-wf-md-setup-protein-cwl. The component BioBB building blocks of the workflow have their own clear and accessible licenses (R1.2). The workflow is associated with detailed provenance about its development, provided as a downloadable RO-Crate object (R1.3).
3.2 Example workflow run
Here, we demonstrate the application of the FAIR principles to a workflow run, which in this case is a dataset that represents an execution of a tissue/tumor prediction workflow for digital pathology using RO-Crate [31].
@@ 3.2 Example workflow run
The workflow run has been assigned a globally unique digital object identifier (F1). Components of the workflow run (e.g. steps, inputs, outputs) are assigned distinct identifiers (F1.1). The workflow run is described using (rich) metadata (F2) that includes the identifier of the workflow run and its unique version (F3). The workflow is not registered, but it is publicly available via GitHub [^6] https://github.com/crs4/deephealth-pipelines, and the workflow run is available on Zenodo (F4) with a DOI.
@@
The workflow run and its components can be retrieved and downloaded using a standardized HTTPS protocol (A1, A1.1, A1.2). The metadata associated with the workflow run and its components are independent from the GitHub code repository and are accessible on Zenodo even if the workflow (or its components) becomes inaccessible (A2).
@@
The workflow run provenance of digital pathology tissue/tumor prediction is generated using CWLProv [32] in the form of RO-Bundle and then converted to an RO-Crate [33] that follows Provenance Run Crate profile [^7] https://www.researchobject.org/workflow-run-crate/profiles/0.1/provenance_run_crate. All these methods represent run information utilizing well-established standards and language (I1). CWLProv and RO-Crate allow vocabularies that are consistent with the FAIR principles (I2). The workflow run provenance is generated using CWLProv and documented as Provenance Run Crate profile showing the exchange of data between components of the workflow. However, the workflow run does not include domain-specific standards for metadata to fulfill I3. The workflow run provenance includes qualified references to other objects and the workflow’s components (I4).
@@
The workflow run is released under MIT License (R1.1) and the associated code is publicly available on GitHub. The component Slaid (a library for applying DL models from the DeepHealth project [^8] https://deephealth-project.eu/ on WSI) of the workflow run is released under MIT License (R1.2). The workflow run contains detailed provenance about the execution, provided as a downloadable RO-Crate object (R1.3). The workflow processes digital pathology images in the formats supported by OpenSlide (R3), a widespread software library for digital pathology images [34].
4 FAIRness beyond the FAIR Principles
@@
Workflows are composites of executable components that need to be validated and tested with clean interfaces, permissive access permissions if remotely executed, and compatible licenses. Work on canonical workflow libraries [37] and building blocks that are interoperable between languages [29], workflow readiness of tools [38], FAIR unit testing [e.g., pytest-workflow for WDL [^9] https://github.com/LUMC/pytest-workflow, test monitoring [e.g. LifeMonitor [^10] https://crs4.github.io/life_monitor/], and benchmarking [e.g. OpenEBench [^11] https://openebench.bsc.es/] are steps towards creating FAIR workflow professional practices that should include peer review, curation, and perhaps even certification.
@@
We should also consider the relationship of FAIR to other services in a workflow’s service ecosystem. Container technologies such as Docker and Singularity play a role in the interoperability, reusability, and reproducibility of computational workflows by providing lightweight, portable, and self-contained environments that include all of the software dependencies and libraries required to run a workflow. In alignment with the FAIR principles, containers can be versioned and shared through registries like Docker Hub, Nexus Repository, or Github Container registry. FAIR workflows should be stored in version-controlled repositories like GitHub and GitLab and indexed by registries like Dockstore [23] and WorkflowHub [22]. Workflow registries like Dockstore and WorkflowHub support findability and accessibility so workflows can be shared, published, and reused as well as downloaded or launched using dedicated infrastructure facilities such as Galaxy Europe. Both support example input and test data components for a workflow specification but do not retain the result components of a workflow run – the provenance and links to datasets that are the resulting products. That is expected to be managed by data repositories like Zenodo and Figshare that in turn need to support FAIR data principles. However, this container approach most probably shows practical limitations when applied to workflows executed on very large-scale, meticulously maintained and pampered HPC environments running on the verge of computational and technical stability. For such cases, reproducibility of workflows might always be an issue.
@@
Workflows are multi-part objects whose data components are as important as their executable steps; all these digital objects should be FAIR too. Community efforts like RO-Crate [39] propose the means to encapsulate workflows and all their components while capturing the context in which they are used (e.g., executions, bibliography, sketches, etc.). An RO-Crate Workflow profile [^12] https://about.workflowhub.eu/Workflow-RO-Crate/ includes references to components which also need to be FAIR. By having workflows as composition of FAIR research outputs, the FAIRness of each contained resource is validated, since the execution of the workflow relies on these resources being available. Services like WorkflowHub (registry), LifeMonitor (test monitoring), and Galaxy (workflow platform) implement RO-Crate for comprehensive workflow representation and exchange as well as a step to sharing metadata and persistent identifiers to implement the FAIR principles of Table 1.
@@
As workflows typically operate on and produce datasets, they are well placed to make those data products FAIR; they provide the process description and automated data provenance collection, and fully fledged WMS can automate processes for generating FAIR data. Nevertheless, workflows have to be well designed to be “FAIR aware”: producing and consuming data in standardized formats, assigning license, handling identifiers, and metadata production, data versioning, etc [12]. Using workflows as automated instruments for “FAIR data by design” can encourage inputs and make outputs to be machine-actionable and more suitable for use in other workflows, for example those that span geographically distinct computing facilities [40].
5 Conclusion

### v2

@@ Abstract
Recent trends within computational and data sciences show an increasing recognition and adoption of computational workflows as tools for productivity and reproducibility that also democratize access to platforms and processing know-how. As digital objects to be shared, discovered, and reused, computational workflows benefit from the FAIR principles, which stand for Findable, Accessible, Interoperable, and Reusable. The Workflows Community Initiative’s FAIR Workflows Working Group (WCI-FW), a global and open community of researchers and developers working with computational workflows across disciplines and domains, has systematically addressed the application of both FAIR data and software principles to computational workflows. We present recommendations with commentary that reflects our discussions and justifies our choices and adaptations. These are offered to workflow users and authors, workflow management system developers, and providers of workflow services as guidelines for adoption and fodder for discussion. The FAIR recommendations for workflows that we propose in this paper will maximize their value as research assets and facilitate their adoption by the wider community.
@@
Computational workflows and why FAIR matters
Scale-ups in research data and the rise of data-driven science have spurred the research community to find solutions for managing and automating complex computational processes in order to ensure efficiency, reproducibility, scalability, collaboration, and quality-assured transparency. Computational workflows are a special kind of software specifically targeted at handling multi-step, multi-code data pipelines, data analyses, and other data-handling operations, especially through the efficient use of computational resources to transform data inputs into desired outputs.
@@
At the simplest end of the complexity spectrum, the workflow language might consist of a script (e.g. Bash, Python, R) or enumerated stages in an electronic research notebook (e.g. Jupyter, RStudio, Apache Zeppelin, etc.) with a set of instructions that use inputs and outputs to pipe the results together. At the other end, workflows might employ a workflow management system (WMS) such as Nextflow [2], Galaxy [3], Snakemake [4], or Parsl [5]. Using a WMS can provide a number of benefits including abstraction, scaling, automation, reproducibility [6], and provenance [7]. A fully featured WMS typically facilitates error handling and restarting, automatic data staging, provenance recording, handling of large datasets and complex computations, task and resource allocation, and distributed task execution to scale over local workstations, cloud resources, or high-performance computing clusters. WMSes and modular containerized software components (e.g. Docker, Singularity) aid portability and reproducibility, but they also face challenges (e.g., security issues when deployed in clusters, steep l earning curve, etc.). In practice, researchers use mixtures of techniques from across the spectrum, such as chaining scripts from a WMS that in turn is launched from a notebook.
@@
A workflow is the formal specification of the data flow and execution control between executable components, the expected datasets, and parameter files. Its complexity can be composed, and its overall process can be abstracted away from its execution. A workflow run is the instantiation of the workflow with inputs (parameters files, input datasets) and outputs (output data, the provenance execution log and lineage of data products).
@@
The use of workflows has accelerated in the past few years. The existence of more than 350 different workflow management systems of varying maturity evidences the increasing popularity of workflows (https://s.apache.org/existing-workflow-systems). Workflows help reduce the burden of manual human effort by automating repetitive and time-consuming tasks. This automation enables efficiency in terms of resource allocation both by allowing humans to return to human-only tasks and by allowing the computational tasks to take advantage of opportunities for automatic optimizations. Workflows also automate critical processes that require standardization and increased reproducibility, improving efficiency by removing humans from the analytical process, at times dispensing with the need for manual management of data flow through a set of processing steps, and reducing the need for repeat analyses due to human error and/or bias. Moreover, automation ensures that computational experiments can be replicated. Workflows also provide a documented record of the computational processes, which helps in understanding, reviewing, and auditing those processes. As scientific activity often includes the exploration of analysis variance, modifying workflows to understand effects and changes on data products is simpler when those workflows are clearly described and comparable. Researchers can reuse workflows to re-analyze and adapt to incorporate new methods, tools, or data sources. Sharing and reusing workflows supports team collaboration and standardizes processes and analysis methods [8].
The Findable, Accessible, Interoperable, and Reusable (FAIR) principles aim to maximize the value and impact of scientific digital objects. Such objects cannot be reused if they cannot be found, accessed, or understood; they cannot be combined if they are not interoperable. The FAIR guidelines for scientific data [9] emphasize making data findable through metadata provision, unique and persistent identifier assignment, ensuring accessibility through open access or controlled access mechanisms, promoting interoperability through standardized formats and metadata schemas, and enabling reusability through clear licensing and documentation. The FAIR principles for research software (FAIR4RS) [10] recommend making software findable through repositories and registries, ensuring accessibility through open licensing and documentation, promoting interoperability through standardization and compatibility with other tools, and enabling reusability through versioning and clear documentation of functionality and dependencies. These principles have made a significant impact on policy and publishing, and they have been taken up by the public and private sector as well as spawning an industry of projects, research programs, and commercial businesses [11, 12].
FAIR Principles applied to computational workflows
@@
The WCI-FW debated for many months whether to define a new, tailored set of principles for computational workflows or to apply the established principles for data and software. We concluded that we would directly apply the principles for FAIR data and software after recognizing that (a) we were naturally developing variations of these established principles because (b) workflows are a hybrid form of software, closely bound to data and even possessing data-like properties [14]. By building on the established FAIR framework, we have maintained consistency and enabled practices for workflows to integrate seamlessly with existing practices for data and software.
@@
  • Do we need to tailor software principles to the specific characteristics of workflows? As workflows are executables, the principles of software should apply. Nevertheless, the characteristics of the heterogeneous components that make up workflows and the expected user behavior of variant reuse, modification, and portability led us to some customization of the principles. Technical information is needed that details each step’s inputs, outputs, dependencies, and computational requirements as well as configuration files, lists of software dependencies, and other information about the operational context. Missing context is one of the main difficulties faced when porting across computing platforms; applying FAIR to workflows requires fully describing the context necessary for executing the workflows as software.
@@
F1. A workflow is assigned a globally unique and persistent identifier. D-F1 and S-F1
@@
F2. A workflow and its components are described with rich metadata. D-F2 and S-F2
F3. Metadata clearly and explicitly include the identifier of the workflow, and workflow versions, that they describe. D-F3 and S-F3
F4. Metadata and workflow are registered or indexed in a searchable FAIR resource. D-F4 and S-F4
A1. Workflow and its components are retrievable by their identifiers using a standardized communications protocol. D-A1 and S-A1
A1.1. The protocol is open, free, and universally implementable. D-A1.1 and S-A1.1
A1.2. The protocol allows for an authentication and authorization procedure, when necessary. D-A1.2 and S-A1.2
A2. Metadata are accessible, even when the workflow is no longer available. D-A2 and S-A2
I1. Workflow and its metadata (including workflow run provenance) use a formal, accessible, shared, transparent, and broadly applicable language for knowledge representation. D-I1 and S-R1.2
@@
I3. Workflow is specified in a way that allows its components to read, write, and exchange data (including intermediate data), in a way that meets domain-relevant standards. D-R3 and S-I1
I4. Workflow and its metadata (including workflow run provenance) include qualified references to other objects and the workflow’s components. D-I3, S-I2, and S-R1.2
R1. Workflow is described with a plurality of accurate and relevant attributes. D-R1 and S-R1
R1.1. Workflow is released with a clear and accessible license. D-R1.1 and S-R1.1
R1.2. Components of the workflow representing levels of granularity are given clear and accessible licenses. D-R1.1 and S-R1.1
R1.3. Workflow is associated with detailed provenance of the workflow and of the products of the workflow. D-R1.2 and S-R1.2
R2. Workflow includes qualified references to other workflows. D-I3 and S-R2
R3. Workflow meets domain-relevant community standards. D-R1.3 and S-R3
@@
Rich metadata for the overall workflows and their components (including inputs and outputs) (F2, F3) should use a formal, accessible language for knowledge representation, such as schema.org (https://schema.org), and standardized metadata such as the Bioschemas schema.org profile for Computational Workflows (https://bioschemas.org/types/ComputationalWorkflow/), used for workflow components in Workflow Run Crate profile [19] and CodeMeta (https://w3id.org/codemeta/v3.0). Other metadata forms for workflows include canonical “Rosetta” languages (e.g. CWL [20], WDL [21]) that abstract from the native workflow language for greater portability and comparability. Such metadata are used for registration and search in dedicated and trusted workflow registries such as WorkflowHub [22] and Dockstore (F4) [23].
Accessibility
@@
Interoperability
The data and software principles combine and blur with principles from Reusability, as we tease apart the FAIR requirements for a workflow, its components (both executable and data), and an instantiated and executed run of the workflow. For example, principle I1 applies to describing the static workflow structure of the specification and capturing dynamic aspects of the run such as execution details and intermediate outputs. The workflow specification plays a role in data in I1, I2, and I4 adapting data principles, as well as the role of metadata of the software, co-opting software principles.
@@
While previously discussing findability, we also highlighted the reusability of workflows as complex compositions intended to be modified, built upon, and incorporated into other workflows, as well as their reusability as runnable executions with modified datasets and parameter files [27]. The data and software principles are adapted to emphasize the composite nature of workflows (both executable and data components) for licensing [28]. Workflows are composed of many parts, potentially from different authors with different intents for how their artifacts may be used; these intents must be clearly stated by the inclusion of licenses that cover all parts of the workflow and sub-parts of the workflows that can be reused as sub-workflows [13].
@@
Examples
Here, we provide concrete examples with detailed descriptions for applying the FAIR principles to both workflow specifications and workflow runs. For each case, we narrate through the FAIR principles sequentially. We recommend the reader to follow along with Table 1 as a guide.
Example workflow specification
The Protein MD Setup workflow sets up a protein molecular dynamic simulation that returns a protein structure and simulated 3D trajectories [29]. It is registered in WorkflowHub (https://doi.org/10.48546/workflowhub.workflow.29.3), where it has been assigned a Digital Object Identifier (DOI) which is globally unique (F1). Different versions of this workflow can be identified using unique digital object identifiers (F1.2). The workflow is described using (rich) metadata (F2) that includes the identifier of the workflow and its unique version (F3). The workflow and associated metadata are registered (F4). Moreover, components of the workflow have their own individual persistent identifiers as the workflow is composed of sub-workflow BioBB building blocks from the BioBB Library (https://workflowhub.eu/projects/11) (F1.2). The workflow and its components can be retrieved and downloaded using a standardized HTTPS protocol (A1, A1.1, A1.2). The metadata associated with the workflow and its components is independent from the GitHub code repository and are accessible on WorkflowHub even if the workflow (or its components) become inaccessible (A2).
@@
The workflow is released under Apache License 2.0 software license (R1.1) and WorkflowHub’s RO-Crate references that the embedded CWL code is archived from the corresponding GitHub source code repository (https://github.com/bioexcel/biobb-wf-md-setup-protein-cwl). The component BioBB building blocks of the workflow have their own clear and accessible licenses (R1.2). The workflow is associated with detailed provenance about its development, provided as a downloadable RO-Crate object (R1.3).
Example workflow run
Here, we demonstrate the application of the FAIR principles to a workflow run. Recall from our previous definitions that a workflow run is the instantiation of a workflow with inputs (parameters files, input datasets) and outputs (output data, the provenance execution log and lineage of data products). In this example, the workflow run is instantiated from a tissue/tumor prediction workflow specification for digital pathology, and it is represented using RO-Crate [31].
@@
The workflow run has been assigned a globally unique DOI (F1), and components of the workflow run (e.g. steps, inputs, outputs) are also assigned their own distinct identifiers (F1.1). The workflow run is described using rich metadata (F2) that includes the identifier of the workflow run and its unique version (F3). The corresponding workflow specification is not registered, but it is publicly available via GitHub (https://github.com/crs4/deephealth-pipelines), and the workflow run is registered on Zenodo, which is a searchable FAIR resource (F4).
@@
The workflow run and its components are also hosted on Zenodo, which means that they can be retrieved and downloaded using the HTTPS protocol, which is standardized (A1), open and free and universally implementable (A1.1), and able to support authentication and authorization procedures (A1.2). The metadata associated with the workflow run and its components are independent from the workflow run and components themselves, and they are accessible on Zenodo even if the workflow run and its components become inaccessible (A2).
@@
Recall from Figure 1 that execution provenance is a product of a workflow run. Here, provenance from the digital pathology tissue/tumor prediction workflow run is generated using CWLProv [32] in the form of RO-Bundle and then converted to an RO-Crate [33] that follows the Provenance Run Crate profile (https://www.researchobject.org/workflow-run-crate/profiles/0.1/provenance_run_crate). All of these methods represent workflow run information by utilizing well-established standards and language (I1). CWLProv and RO-Crate allow vocabularies that are consistent with the FAIR principles (I2). The workflow run provenance is generated using CWLProv and documented as a Provenance Run Crate profile showing the exchange of data between components of the workflow. The workflow run does not, however, include domain-specific standards for metadata to fulfill (which means it does not satisfy I3). The workflow run provenance does include qualified references to other objects and the workflow’s components (I4).
@@
The workflow run is described with many accurate and relevant attributes (R1), and it is released under the clear and accessible MIT License (R1.1). One of the workflow run’s components, Slaid, is a library for applying DL models from the DeepHealth project (https://deephealth-project.eu/) on whole slide images (WSI) that is also released independently under the MIT License (R1.2). The workflow run contains detailed provenance about the execution, provided as a downloadable RO-Crate object (R1.3). The workflow run processes digital pathology images in the formats supported by OpenSlide (R3), which is a widespread software library that represents a community standard for digital pathology images [34].
FAIRness beyond the FAIR Principles
@@
Workflows are composites of executable components that need to be validated and tested with clean interfaces, permissive access permissions if remotely executed, and compatible licenses. Work on canonical workflow libraries [37] and building blocks that are interoperable between languages [29], workflow readiness of tools [38], FAIR unit testing like pytest-workflow for WDL (https://github.com/LUMC/pytest-workflow), test monitoring like LifeMonitor(https://crs4.github.io/life_monitor/), and benchmarking like OpenEBench (https://openebench.bsc.es/) are steps towards creating FAIR workflow professional practices that should include peer review, curation, and perhaps even certification.
@@
We should also consider the relationship of FAIR to other services in a workflow’s service ecosystem. Container technologies such as Docker and Singularity play a role in the interoperability, reusability, and reproducibility of computational workflows by providing lightweight, portable, and self-contained environments that include all of the software dependencies and libraries required to run a workflow. In alignment with the FAIR principles, containers can be versioned and shared through registries like Docker Hub, Nexus Repository, or Github Container registry, but these registries are not truly FAIR because old versions may not be preserved. FAIR workflows should be stored in version-controlled repositories like GitHub and GitLab and indexed by registries like Dockstore [23] and WorkflowHub [22]. Workflow registries like Dockstore and WorkflowHub support findability and accessibility so workflows can be shared, published, and reused as well as downloaded or launched using dedicated infrastructure facilities such as Galaxy Europe. Both support example input and test data components for a workflow specification but do not retain the result components of a workflow run – the provenance and links to datasets that are the resulting products. That is expected to be managed by data repositories like Zenodo and Figshare that in turn need to support FAIR data principles. However, this container approach most probably shows practical limitations when applied to workflows executed on very large-scale, meticulously maintained and pampered HPC environments running on the verge of computational and technical stability. For such cases, reproducibility of workflows might always be an issue.
@@
Workflows are multi-part objects whose data components are as important as their executable steps; all these digital objects should be FAIR too. Community efforts like RO-Crate [39] propose the means to encapsulate workflows and all their components while capturing the context in which they are used (e.g., executions, bibliography, sketches, etc.). An RO-Crate Workflow profile (https://about.workflowhub.eu/Workflow-RO-Crate/) includes references to components which also need to be FAIR. By having workflows as composition of FAIR research outputs, the FAIRness of each contained resource is validated, since the execution of the workflow relies on these resources being available. Services like WorkflowHub (registry), LifeMonitor (test monitoring), and Galaxy (workflow platform) implement RO-Crate for comprehensive workflow representation and exchange as well as a step to sharing metadata and persistent identifiers to implement the FAIR principles of Table 1.
@@
As workflows typically operate on and produce datasets, they are well placed to make those data products FAIR; they provide the process description and automated data provenance collection, and a fully fledged WMS can automate processes for generating FAIR data. Nevertheless, workflows have to be well designed to be “FAIR aware”: producing and consuming data in standardized formats, assigning license, handling identifiers, and metadata production, data versioning, etc [14]. Using workflows as automated instruments for “FAIR data by design” can encourage inputs and make outputs to be machine-actionable and more suitable for use in other workflows, for example those that span geographically distinct computing facilities [40].
Conclusion